# Nonsmooth Implicit Differentiation: Deterministic and Stochastic Convergence Rates

### v1

@@ Appendix D Stochastic Implicit Differentiation
$=3\lVert y\rVert^{2}\bigg{(}\sigma_{\lambda}(k)+\frac{1}{(1-q)^{4}}\Big{(}C^{2}_{G,\lambda}\frac{\operatorname{Var}[\hat{T}(z_{t},\xi)\,|\,w_{t}(\lambda),y]}{J}+(C_{G,\lambda}+C_{T,\lambda})^{2}\Delta_{t}^{2}\Big{)}\bigg{)}$
$\leq 3\lVert y\rVert^{2}\bigg{(}\sigma_{\lambda}(k)+\frac{1}{(1-q)^{4}}\Big{(}C^{2}_{G,\lambda}\frac{B_{3}(\Delta_{t}^{2})}{J}+(C_{G,\lambda}+C_{T,\lambda})^{2}\Delta_{t}^{2}\Big{)}\bigg{)}$
@@
$(2)\leq 6\lVert D_{\Phi,2}(z)\rVert^{2}_{\sup}\lVert y\rVert^{2}\bigg{(}\sigma_{\lambda}(k)+\frac{1}{(1-q)^{4}}\Big{(}C^{2}_{G,\lambda}\frac{B_{3}(\Delta_{t}^{2})}{J}+(C_{G,\lambda}+C_{T,\lambda})^{2}\Delta_{t}^{2}\Big{)}\bigg{)}.$
@@
$\leq 12\lVert y\rVert^{2}\big{(}\sigma_{\lambda}(k)+\kappa^{2}\big{)}\bigg{(}\frac{\sigma_{2}^{\prime}}{J}+[C_{T,\lambda}+C_{G,\lambda}M_{T,\lambda}]^{2}\Delta_{t}^{2}+C^{2}_{G,\lambda}M_{T,\lambda}^{2}\frac{\sigma_{1}+3\sigma_{2}(2\Delta_{t}^{2}+\Delta_{\lambda}^{2})}{J}\bigg{)}$
$\qquad+6\lVert y\rVert^{2}\lVert D_{\Phi,2}(w(\lambda),\lambda)\rVert^{2}_{\sup}\bigg{(}\sigma_{\lambda}(k)+\kappa^{4}\Big{(}C^{2}_{G,\lambda}\frac{\sigma_{1}+3\sigma_{2}(2\Delta_{t}^{2}+\Delta_{\lambda}^{2})}{J}+(C_{G,\lambda}+C_{T,\lambda})^{2}\Delta_{t}^{2}\Big{)}\bigg{)},$
where in the last equation we used the expression for $B_{3}(\Delta_{t}^{2})$ and $\kappa=(1-q)^{-1}$. Taking the expectation $\mathbb{E}[\cdot\,|\,w_{t}(\lambda)]$ and recalling the hypothesis on $\lVert y\rVert^{2}$ and $\Delta^{2}_{t}$ in 5.3(i)(iii), the statement follows. ∎
@@
[…]
for some $\hat{\sigma}_{1},\hat{\sigma}_{2}>0$. Let $\eta_{i}=\beta/(\gamma+i)$, with $\beta>1/(1-q^{2})$ and $\gamma\geq\beta(1+\hat{\sigma}_{2}^{2})$. Let $(\xi_{i})_{i\in\mathbb{N}}$ be a sequence of i.i.d copies of $\xi$ and let $(v_{i})_{i\in\mathbb{N}}$ be such that $v_{0}=0$ and for $i=0,1,\dots$
@@
1:  Input: $k\in\mathbb{N}$, $u_{1},u_{2},y\in\mathbb{R}^{d}$, $\hat{T}$, $G$, $\bm{\xi}^{(2)}=(\xi^{(2)}_{i})_{1\leq i\leq k}$.
@@
$\operatorname{Var}[\hat{\Psi}(v,\xi)]\leq\operatorname{Var}[\partial_{1}\hat{T}(u_{1},\lambda,\xi)]\lVert v\rVert^{2}.$
@@
$\operatorname{Var}[\hat{\Psi}(v,\xi)]\leq\frac{2}{(1-q)^{2}}\operatorname{Var}[\partial_{1}\hat{T}(u_{1},\lambda,\xi)]\big{(}\lVert v-\Psi(v)\rVert^{2}+\lVert y\rVert^{2}\big{)}\leq\frac{2\sigma_{1}^{\prime}}{(1-q)^{2}}\lVert\Psi(v)-v\rVert^{2}+\frac{2\sigma_{1}^{\prime}\lVert y\rVert^{2}}{(1-q)^{2}}.$
Therefore, the first part of the statement follows from Lemma D.2 and from $\lVert\bar{v}\rVert\leq\lVert y\rVert(1-q)^{-1}$ (a consequence of (28)). The last part follows by (27) and the fact that
@@
By applying Lemma D.3 with $u_{1}=w_{t}(\lambda)$ and $u_{2}=\bar{T}(w_{t}(\lambda),\lambda)$ we obtain that 5.3(ii) (the rate on the mean square error of $v_{k}$) is satisfied with $\sigma_{\lambda}(k)=O(\kappa^{5}k^{-1})$. The statement follows by applying Theorem 5.4 and substituting the rates $\rho_{\lambda}(t)$ and $\sigma_{\lambda}(k)$. ∎
@@ Appendix E Bilevel Optimization
In this section we consider Problem Equation 20 and we make the following assumption.
@@
Note that similarly to $\Phi$, since $E$ satisfies Assumption E.1, an application of Lemma 2.5 to the map $E$ yields
@@
Under E.1, for every $\lambda\in\Lambda$, there exist $R_{E,\lambda}>0$ such that for every $u\in\mathbb{R}^{d}$ and $k\in\{1,2\}$
$\operatorname{\it{e}}(D_{E,k}(u,\lambda),D_{E,k}(w(\lambda),\lambda))\leq C_{E,\lambda}\lVert u-w(\lambda)\rVert$
@@ E.1 Deterministic Case
$\operatorname{\it{e}}(D_{f_{t}}^{k}(z_{t}),D_{f}^{\operatorname{fix}}(z))=O(\kappa e^{-k/\kappa}+\kappa^{2}\Delta_{t})$
@@
$\operatorname{\it{e}}(D_{f_{t}}(z_{t}),D_{f}^{\operatorname{fix}}(z))=O(\kappa te^{-\kappa/t}).$
@@
Recalling the expressions for $D_{f_{t}}$ and $D_{f}^{\operatorname{fix}}$ we obtain
$\operatorname{\it{e}}(D_{f_{t}}(z_{t}),D_{f}^{\operatorname{fix}}(z))$ $\leq\lVert D_{E,1}(z_{t})\rVert_{\sup}\operatorname{\it{e}}(D_{w_{t}}(\lambda),D_{w}^{\operatorname{fix}}(\lambda))$
$\quad+\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}\operatorname{\it{e}}(D_{E,1}(z_{t}),D_{E,1}(z))$
$\quad+\operatorname{\it{e}}(D_{E,2}(z_{t}),D_{E,2}(z))$
$\leq(\lVert D_{E,1}(z)\rVert_{\sup}+C_{E,\lambda}\Delta_{t})\times O(\kappa te^{-t/\kappa})$
$+\Big{(}\frac{\lVert D_{\Phi,2}(z_{t})\rVert_{\sup}}{1-q}+1\Big{)}C_{E,\lambda}\Delta_{t}$
$\leq(\lVert D_{E,1}(z)\rVert_{\sup}+C_{E,\lambda}\Delta_{0})\times\underbrace{O(\kappa te^{-t/\kappa})}_{(*)}$
$+\Big{(}\frac{\lVert D_{\Phi,2}(z)\rVert_{\sup}+C_{\Phi,\lambda}\Delta_{0}}{1-q}+1\Big{)}C_{E,\lambda}\Delta_{0}e^{-t/\kappa},$
where we used the ITD bound in Theorem 4.1 and the Lemma E.2. A very similar proof can be done for AID-FP by changing the (*)(*)( * ) term to $O(\kappa e^{-k/\kappa}+\kappa^{2}e^{-t/\kappa}).$
@@ E.2 Stochastic Case
$E(w,\lambda)=\mathbb{E}[\hat{E}(w,\lambda,\zeta)],\ \Phi(w,\lambda)=G(\mathbb{E}[\hat{T}(w,\lambda,\xi)],\lambda).$
@@
The assumption above is verified e.g. for the logistic and for the cross-entropy loss. Moreover, the assumptions on $\hat{E}$ are the following.
@@
    $\hat{E}\colon\mathbb{R}^{d}\times O_{\Lambda}\times\mathcal{Z}\to\mathbb{R}^{d}$, $\zeta$ is a random variable with value in $\mathcal{Z}$ and $\mathbb{E}[\hat{E}(u,\lambda,\zeta)]=E(u,\lambda)$.
@@
    For every $z\in\mathcal{Z}$, $\hat{E}(\cdot,\cdot,z)$ is path differentiable and
    $\partial_{1}\hat{E}\colon\mathbb{R}^{d}\times O_{\Lambda}\times\mathcal{Z}\to\mathbb{R}^{d},\ \ \partial_{1}\hat{E}\colon\mathbb{R}^{d}\times O_{\Lambda}\times\mathcal{Z}\to\mathbb{R}^{d}$
    are selections of $D_{\hat{E},1}$ and $D_{\hat{E},2}$ respectively and there exist $\sigma_{E,1},\sigma_{E,2}\geq 0$ such that for every $u\in\mathbb{R}^{d}$ and $\lambda\in\Lambda$
    $\mathbb{E}[\partial_{1}\hat{E}(u,\lambda,\zeta)]=\partial_{1}E(u,\lambda)\in D_{E,1}(u,\lambda),\quad\mathbb{E}[\partial_{2}\hat{E}(u,\lambda,\zeta)]=\partial_{2}E(u,\lambda)\in D_{E,2}(u,\lambda),$
    $\operatorname{Var}[\partial_{1}\hat{E}(u,\lambda,\zeta)]\leq\sigma_{E,1},\quad\operatorname{Var}[\partial_{2}\hat{E}(u,\lambda,\zeta)]\leq\sigma_{E,2}.$
@@
Let 5.1, 5.2, E.1, E.4, E.5 hold and let $\kappa=(1-q)^{-1}$. Also, suppose that $\mathbb{E}[\lVert w_{t}(\lambda)-w(\lambda)\rVert]\leq\rho_{\lambda}(t)$, for every $t\in\mathbb{N}$. Then the output ^$\hat{\nabla}f(\lambda)$ of NSID-Bilevel (Algorithm 2) where NSID uses step sizes $\eta_{i}=O(i^{-1})$ satisfies
@@
Therefore, by setting e.g. $t=k=J_{1}=J_{2}$ we have
@@
We recall that
$D_{f}^{\operatorname{imp}}(\lambda)$ ${:=}D_{w}^{\operatorname{imp}}(\lambda)^{\top}\,D_{E,1}(w(\lambda),\lambda)+D_{E,2}(w(\lambda),\lambda)$
^$\hat{\nabla}f(\lambda)$ $=r(z_{t})+\partial_{2}\bar{E}(z_{t}).$
Then
$\operatorname{\it{e}}($$\hat{\nabla}f(\lambda),D_{f}^{\operatorname{imp}}(\lambda))$
$\leq\operatorname{\it{e}}(r(z_{t}),D_{w}^{\operatorname{imp}}(\lambda)^{\top}\,D_{E,1}(w(\lambda),\lambda))+\operatorname{\it{e}}(\partial_{2}\bar{E}(z_{t}),D_{E,2}(w(\lambda),\lambda))$
$\leq\operatorname{\it{e}}(r(z_{t}),D_{w}^{\operatorname{imp}}(\lambda)^{\top}\,\partial_{1}\bar{E}(z_{t}))+\operatorname{\it{e}}(D_{w}^{\operatorname{imp}}(\lambda)^{\top}\,\partial_{1}\bar{E}(z_{t}),D_{w}^{\operatorname{imp}}(\lambda)^{\top}\,D_{E,1}(w(\lambda),\lambda))$
$\qquad+\lVert\partial_{2}\bar{E}(z_{t})-\partial_{2}E(z_{t})\rVert+\operatorname{\it{e}}(\partial_{2}E(z_{t}),D_{E,2}(w(\lambda),\lambda))$
$\leq\operatorname{\it{e}}(r(z_{t}),D_{w}^{\operatorname{imp}}(\lambda)^{\top}\,\partial_{1}\bar{E}(z_{t}))+\lVert D_{w}^{\operatorname{imp}}(\lambda)\rVert_{\sup}\big{(}\lVert\partial_{1}\bar{E}(z_{t})-\partial_{1}E(z_{t})\rVert+\operatorname{\it{e}}(\partial_{1}E(z_{t}),D_{E,1}(w(\lambda),\lambda))\big{)}$
$\qquad+\lVert\partial_{2}\bar{E}(z_{t})-\partial_{2}E(z_{t})\rVert+C_{E,\lambda}\lVert w_{t}(\lambda)-w(\lambda)\rVert$
$\leq\operatorname{\it{e}}(r(z_{t}),D_{w}^{\operatorname{imp}}(\lambda)^{\top}\,\partial_{1}\bar{E}(z_{t}))+\lVert D_{w}^{\operatorname{imp}}(\lambda)\rVert_{\sup}\big{(}\lVert\partial_{1}\bar{E}(z_{t})-\partial_{1}E(z_{t})\rVert+C_{E,\lambda}\lVert w_{t}(\lambda)-w(\lambda)\rVert\big{)}$
$\qquad+\lVert\partial_{2}\bar{E}(z_{t})-\partial_{2}E(z_{t})\rVert+C_{E,\lambda}\lVert w_{t}(\lambda)-w(\lambda)\rVert$
$\leq\operatorname{\it{e}}(r(z_{t}),D_{w}^{\operatorname{imp}}(\lambda)^{\top}\,\partial_{1}\bar{E}(z_{t}))+C_{E,\lambda}(1+\lVert D_{w}^{\operatorname{imp}}(\lambda)\rVert_{\sup})\Delta_{t}$
$\qquad+\lVert\partial_{2}\bar{E}(z_{t})-\partial_{2}E(z_{t})\rVert+\lVert D_{w}^{\operatorname{imp}}(\lambda)\rVert_{\sup}\lVert\partial_{1}\bar{E}(z_{t})-\partial_{1}E(z_{t})\rVert$
@@
$\mathbb{E}[\operatorname{\it{e}}(\hat{\nabla}f(\lambda),$ $D_{f}^{\operatorname{imp}}(\lambda))^{2}\,|\,w_{t}(\lambda)]$
$\leq 4\bigg{(}\mathbb{E}[\operatorname{\it{e}}(r(z_{t}),D_{w}^{\operatorname{imp}}(\lambda)^{\top}\,\partial_{1}\bar{E}(z_{t}))^{2}\,|\,w_{t}(\lambda)]+C^{2}_{E,\lambda}(1+\lVert D_{w}^{\operatorname{imp}}(\lambda)\rVert_{\sup})^{2}\Delta^{2}_{t}$
$\qquad\qquad+\frac{\operatorname{Var}[\partial_{2}\hat{E}(z_{t},\zeta)\,|\,w_{t}(\lambda)]}{J_{1}}+\lVert D_{w}^{\operatorname{imp}}(\lambda)\rVert^{2}_{\sup}\frac{\operatorname{Var}[\partial_{1}\hat{E}(z_{t},\zeta)\,|\,w_{t}(\lambda)]}{J_{1}}\bigg{)}$
$\leq 4\bigg{(}\mathbb{E}[\operatorname{\it{e}}(r(z_{t}),D_{w}^{\operatorname{imp}}(\lambda)^{\top}\,\partial_{1}\bar{E}(z_{t}))^{2}\,|\,w_{t}(\lambda)]+C^{2}_{E,\lambda}\bigg{(}1+\frac{\lVert D_{\Phi,2}(w(\lambda),\lambda)\rVert_{\sup}}{1-q}\bigg{)}^{2}\Delta^{2}_{t}$
$\qquad\qquad+\frac{\sigma_{E,2}}{J_{1}}+\frac{\lVert D_{\Phi,2}(w(\lambda),\lambda)\rVert^{2}_{\sup}}{(1-q)^{2}}\frac{\sigma_{E,1}}{J_{1}}\bigg{)}.$ (29)
Now, we note that
$\mathbb{E}\big{[}\lVert\partial_{1}\bar{E}(z_{t})\rVert^{2}\,|\,w_{t}(\lambda)\big{]}\leq 2\mathbb{E}\big{[}\lVert\partial_{1}\bar{E}(z_{t})-\partial_{1}E(z_{t})\rVert^{2}\,|\,w_{t}(\lambda)\big{]}+2\lVert D_{E,1}(z_{t})\rVert^{2}_{\sup}\leq\frac{2\sigma_{E,1}}{J_{1}}+2B_{E,\lambda}.$ (30)
Therefore, taking the total expectation in (E.2) and recalling Theorem 5.4, we get

### v2

@@ Appendix D Stochastic Implicit Differentiation
$=3\lVert y\rVert^{2}\bigg{(}\sigma_{\lambda}(k)+\frac{1}{(1-q)^{4}}\Big{(}C^{2}_{G,\lambda}\frac{\operatorname{Var}[\hat{T}_{\xi}(z_{t})\,|\,w_{t}(\lambda),y]}{J}+(C_{G,\lambda}+C_{T,\lambda})^{2}\Delta_{t}^{2}\Big{)}\bigg{)}$
$\leq 3\lVert y\rVert^{2}\bigg{(}\sigma_{\lambda}(k)+\frac{1}{(1-q)^{4}}\Big{(}C^{2}_{G,\lambda}\frac{b_{\lambda}(\Delta_{t}^{2})}{J}+(C_{G,\lambda}+C_{T,\lambda})^{2}\Delta_{t}^{2}\Big{)}\bigg{)}.$
@@
$\mathbb{E}[(2)\,|\,w_{t}(\lambda),y]\leq 6\lVert D_{\Phi,2}(w(\lambda),\lambda)\rVert^{2}_{\sup}\lVert y\rVert^{2}\bigg{(}\sigma_{\lambda}(k)+\frac{1}{(1-q)^{4}}\Big{(}C^{2}_{G,\lambda}\frac{b_{\lambda}(\Delta_{t}^{2})}{J}+(C_{G,\lambda}+C_{T,\lambda})^{2}\Delta_{t}^{2}\Big{)}\bigg{)}.$
@@
$\leq 12\lVert y\rVert^{2}\big{(}\sigma_{\lambda}(k)+\kappa^{2}\big{)}\bigg{(}\frac{\sigma_{2}^{\prime}}{J}+[C_{T,\lambda}+C_{G,\lambda}M_{T,\lambda}]^{2}\Delta_{t}^{2}+C^{2}_{G,\lambda}M_{T,\lambda}^{2}\frac{\sigma_{1}+3\sigma_{2}(2\Delta_{t}^{2}+a_{\lambda}^{2})}{J}\bigg{)}$
$\qquad+6\lVert y\rVert^{2}\lVert D_{\Phi,2}(w(\lambda),\lambda)\rVert^{2}_{\sup}\bigg{(}\sigma_{\lambda}(k)+\kappa^{4}\Big{(}C^{2}_{G,\lambda}\frac{\sigma_{1}+3\sigma_{2}(2\Delta_{t}^{2}+a_{\lambda}^{2})}{J}+(C_{G,\lambda}+C_{T,\lambda})^{2}\Delta_{t}^{2}\Big{)}\bigg{)},$
where we used the expression for $b_{\lambda}(\Delta_{t}^{2})$ and $\kappa=(1-q)^{-1}$. Taking the expectation $\mathbb{E}[\cdot\,|\,w_{t}(\lambda)]$ and recalling the hypothesis on $\lVert y\rVert^{2}$ and $\Delta^{2}_{t}$ in 5.4(i)(iii), the statement follows. ∎
@@
$\mathbb{E}[\hat{\Psi}(v,\xi)]=\Psi(v)\quad\text{and}\quad\operatorname{Var}[\hat{\Psi}(v,\xi)]\leq\hat{\sigma}_{1}+\hat{\sigma}_{2}\lVert\Psi(v)-v\rVert^{2},$
for some $\hat{\sigma}_{1},\hat{\sigma}_{2}>0$. Let $\eta_{i}=\beta/(\gamma+i)$, with $\beta>1/(1-q^{2})$ and $\gamma\geq\beta(1+\hat{\sigma}_{2})$. Let $(\xi_{i})_{i\in\mathbb{N}}$ be a sequence of i.i.d copies of $\xi$ and let $(v_{i})_{i\in\mathbb{N}}$ be such that $v_{0}=0$ and for $i=0,1,\dots$
@@
1:  Input: $k\in\mathbb{N}$, $u_{1},u_{2},y\in\mathbb{R}^{d}$, $\bm{\xi}^{(2)}=(\xi^{(2)}_{i})_{1\leq i\leq k}$.
@@
$\operatorname{Var}[\hat{\Psi}(v,\xi)]\leq\operatorname{Var}[\partial_{1}\hat{T}_{\xi}(u_{1},\lambda)]\lVert v\rVert^{2}.$
@@
$\operatorname{Var}[\hat{\Psi}(v,\xi)]\leq\frac{2}{(1-q)^{2}}\operatorname{Var}[\partial_{1}\hat{T}_{\xi}(u_{1},\lambda)]\big{(}\lVert v-\Psi(v)\rVert^{2}+\lVert y\rVert^{2}\big{)}\leq\frac{2\sigma_{1}^{\prime}}{(1-q)^{2}}\lVert\Psi(v)-v\rVert^{2}+\frac{2\sigma_{1}^{\prime}\lVert y\rVert^{2}}{(1-q)^{2}}.$
Therefore, the first part of the statement follows from Lemma D.2 and from $\lVert\bar{v}\rVert\leq\lVert y\rVert(1-q)^{-1}$ (a consequence of (30)). The last part follows by (29), the equations
@@
By applying Lemma D.3 with $u_{1}=w_{t}(\lambda)$ and $u_{2}=\bar{T}(w_{t}(\lambda),\lambda)$ we obtain that 5.4(ii) (the rate on the mean square error of $v_{k}$) is satisfied with $\sigma_{\lambda}(k)=O(\kappa^{5}k^{-1})$. The statement follows by applying Theorem 5.5 and substituting the rates $\rho_{\lambda}(t)$ and $\sigma_{\lambda}(k)$. ∎
@@ Appendix E Bilevel Optimization
In this section we consider Problem (22) and we make the following assumption.
@@
Note that similarly to $\Phi$, since $E$ satisfies Assumption E.1, a direct application of Lemma 2.5 to the map $E$ yields
@@
Under E.1, for every $\lambda\in\Lambda$, there exist $R_{E,\lambda}>0$ such that for every $u\in\mathbb{R}^{d}$
$\operatorname{\it{e}}(D_{E}(u,\lambda),D_{E}(w(\lambda),\lambda))\leq C_{E,\lambda}\lVert u-w(\lambda)\rVert,$
@@ E.1 Deterministic Case
$\operatorname{\it{e}}(D_{f_{t}}^{k}(\lambda),D_{f}^{\operatorname{fix}}(\lambda))=O(\kappa e^{-k/\kappa}+\kappa^{2}\Delta_{t})$
@@
$\operatorname{\it{e}}(D_{f_{t}}(\lambda),D_{f}^{\operatorname{fix}}(\lambda))=O(\kappa te^{-\kappa/t}).$
@@
For simplicity, let $\mathcal{A}=D_{E}(w(\lambda),\lambda)$, $\mathcal{A}_{t}=D_{E}(w_{t}(\lambda),\lambda)$ and recall that
$D_{f_{t}}(\lambda)=\mathcal{A}_{t}(D_{w_{t}}(\lambda)),\qquad D_{f}^{\operatorname{fix}}(\lambda)=\mathcal{A}(D_{w}^{\operatorname{fix}}(\lambda)).$
Using the properties of excess in Lemma B.1 we obtain, for BITD:
$\operatorname{\it{e}}(D_{f_{t}}(\lambda),D_{f}^{\operatorname{fix}}(\lambda))$ $\leq\operatorname{\it{e}}(\mathcal{A}_{t}(D_{w_{t}}(\lambda)),\mathcal{A}_{t}(D_{w}^{\operatorname{fix}}(\lambda)))+\operatorname{\it{e}}(\mathcal{A}_{t}(D_{w}^{\operatorname{fix}}(\lambda)),\mathcal{A}(D_{w}^{\operatorname{fix}}(\lambda)))$
$\leq\lVert D_{E,1}(w_{t}(\lambda),\lambda)\rVert_{\sup}\operatorname{\it{e}}(D_{w_{t}}(\lambda),D_{w}^{\operatorname{fix}}(\lambda))$
$\quad+\left(1+\lVert D_{w}^{\operatorname{fix}}(\lambda)\rVert_{\sup}\right)\operatorname{\it{e}}(D_{E}(w_{t}(\lambda),\lambda),D_{E}(w(\lambda),\lambda))$
$\leq(\lVert D_{E,1}(w(\lambda),\lambda)\rVert_{\sup}+C_{E,\lambda}\Delta_{t})\operatorname{\it{e}}(D_{w_{t}}(\lambda),D_{w}^{\operatorname{fix}}(\lambda))$
$+\Big{(}\frac{\lVert D_{\Phi,2}(w(\lambda),\lambda)\rVert_{\sup}}{1-q}+1\Big{)}C_{E,\lambda}\Delta_{t}$
$\leq(\lVert D_{E,1}(w(\lambda),\lambda)\rVert_{\sup}+C_{E,\lambda}\Delta_{0})\times\underbrace{O(\kappa te^{-t/\kappa})}_{(*)}$
$+\Big{(}\frac{\lVert D_{\Phi,2}(w(\lambda),\lambda)\rVert_{\sup}}{1-q}+1\Big{)}C_{E,\lambda}\Delta_{0}e^{-t/\kappa},$
where we used $\Delta_{t}\leq\Delta_{0}e^{-t/\kappa}<\Delta_{0}$, the ITD bound in Theorem 4.1 and Lemma E.2. A very similar proof can be done for AID-FP by changing the ()(*)( ∗ ) term to $O(\kappa e^{-k/\kappa}+\kappa^{2}e^{-t/\kappa}).$
@@ E.2 Stochastic Case
$E(w,\lambda)=\mathbb{E}[\hat{E}_{\zeta}(w,\lambda)],\ \Phi(w,\lambda)=G(\mathbb{E}[\hat{T}_{\xi}(w,\lambda)],\lambda).$
@@
The assumption above is verified e.g., for the logistic and for the cross-entropy loss. Moreover, the assumptions on $\hat{E}$ are the following.
@@
$\zeta$ is a random variable with values in $\mathcal{Z}$ and for every $z\in\mathcal{Z}$
@@
    $\hat{E}_{z}\colon\mathbb{R}^{d}\times O_{\Lambda}\to\mathbb{R}^{d}$, $\mathbb{E}[\hat{E}_{\zeta}(u,\lambda)]=E(u,\lambda)$.
@@
    $\hat{E}_{z}$ is path differentiable with conservative derivative $D_{\hat{E}}$ and $E_{z}^{\prime}$ is a selection of $D_{\hat{E}}$ such that $\hat{E}_{z}^{\prime}(u,\lambda)=[\partial_{1}\hat{E}_{z}(u,\lambda),\partial_{2}\hat{E}_{\zeta}(u,\lambda)]$ and there exist $\sigma_{E,1},\sigma_{E,2}\geq 0$ such that for every $u\in\mathbb{R}^{d}$ and $\lambda\in\Lambda$
    $\mathbb{E}\big{[}E_{\zeta}^{\prime}(u,\lambda)\big{]}=E^{\prime}(u,\lambda)\in D_{E}(u,\lambda),\quad\operatorname{Var}[\partial_{1}\hat{E}_{\zeta}(u,\lambda)]\leq\sigma_{E,1},\quad\operatorname{Var}[\partial_{2}\hat{E}_{\zeta}(u,\lambda)]\leq\sigma_{E,2}.$
@@
Let 5.1, 5.2, E.1, E.4, E.5 hold and let $\kappa=(1-q)^{-1}$. Also, suppose that $\mathbb{E}[\lVert w_{t}(\lambda)-w(\lambda)\rVert]\leq\rho_{\lambda}(t)$, for every $t\in\mathbb{N}$. Then the output ^$\hat{\nabla}f(\lambda)$ of NSID-Bilevel (Algorithm 2) where NSID uses step sizes $\eta_{i}=\Theta(i^{-1})$ satisfies
@@
Therefore, by setting e.g., $t=k=J_{1}=J_{2}$ we have
@@
For simplicity, let $z_{t}=(w_{t}(\lambda),\lambda)$, $z=(w(\lambda),\lambda)$, $\mathcal{A}=D_{E}(w(\lambda),\lambda)$, $\mathcal{B}_{t}=\{\bar{E}^{\prime}(z_{t})\}$. We also recall that
$D_{f}^{\operatorname{imp}}(\lambda):=\mathcal{A}(D_{w}^{\operatorname{imp}}(\lambda)),\qquad\hat{\nabla}f(\lambda)^{\top}=r(z_{t})^{\top}+\partial_{2}\bar{E}(z_{t}),$
with $r(z_{t})$ which is an estimator of $D_{w}^{\operatorname{imp}}(\lambda)^{\top}\partial_{1}\bar{E}(z_{t})$. Then, using the properties in Lemma B.1 and noting that $\mathcal{B}_{t}(D_{w}^{\operatorname{imp}}(\lambda))=\partial_{1}\bar{E}(z_{t})D_{w}^{\operatorname{imp}}(\lambda)+\partial_{2}\bar{E}(z_{t})$, we have
$\operatorname{\it{e}}($$\hat{\nabla}f(\lambda)^{\top},D_{f}^{\operatorname{imp}}(\lambda))$
$\leq\operatorname{\it{e}}(\hat{\nabla}f(\lambda)^{\top},\mathcal{B}_{t}(D_{w}^{\operatorname{imp}}(\lambda))+\operatorname{\it{e}}(\mathcal{B}_{t}(D_{w}^{\operatorname{imp}}(\lambda)),\mathcal{A}(D_{w}^{\operatorname{imp}}(\lambda)))$
$\leq\operatorname{\it{e}}(r(z_{t}),D_{w}^{\operatorname{imp}}(\lambda)^{\top}\,\partial_{1}\bar{E}(z_{t}))+(1+\lVert D_{w}^{\operatorname{imp}}(\lambda)\rVert_{\sup})\operatorname{\it{e}}(\bar{E}^{\prime}(z_{t}),D_{E}(z))$
$\leq\operatorname{\it{e}}(r(z_{t}),D_{w}^{\operatorname{imp}}(\lambda)^{\top}\,\partial_{1}\bar{E}(z_{t}))+(1+\lVert D_{w}^{\operatorname{imp}}(\lambda)\rVert_{\sup})(\lVert\bar{E}^{\prime}(z_{t})-E^{\prime}(z_{t})\rVert+\operatorname{\it{e}}(E^{\prime}(z_{t}),D_{E}(z)))$
$\leq\operatorname{\it{e}}(r(z_{t}),D_{w}^{\operatorname{imp}}(\lambda)^{\top}\,\partial_{1}\bar{E}(z_{t}))+(1+\lVert D_{w}^{\operatorname{imp}}(\lambda)\rVert_{\sup})(\lVert\bar{E}^{\prime}(z_{t})-E^{\prime}(z_{t})\rVert+C_{E,\lambda}\Delta_{t})$
Moreover, let $\tilde{\mathbb{E}}=\mathbb{E}[\,\cdot\,|\,w_{t}(\lambda)]$, we have that
$\tilde{\mathbb{E}}[\lVert\bar{E}^{\prime}(z_{t})-E^{\prime}(z_{t})\rVert^{2}]$ $=\tilde{\mathbb{E}}[\lVert\partial_{1}\bar{E}(z_{t})-\partial_{1}E(z_{t})\rVert^{2}]+\tilde{\mathbb{E}}[\lVert\partial_{2}\bar{E}(z_{t})-\partial_{2}E(z_{t})\rVert^{2}]$
$\leq\frac{\operatorname{Var}[\partial_{1}\hat{E}_{\zeta}(z_{t})\,|\,w_{t}(\lambda)]+\operatorname{Var}[\partial_{2}\hat{E}_{\zeta}(z_{t})\,|\,w_{t}(\lambda)]}{J_{1}}\leq\frac{\sigma_{E,1}+\sigma_{E,2}}{J_{1}}.$
@@
$\tilde{\mathbb{E}}[$ $\operatorname{\it{e}}(\hat{\nabla}f(\lambda)^{\top},D_{f}^{\operatorname{imp}}(\lambda))^{2}]$
$\leq 3\left(\tilde{\mathbb{E}}[\operatorname{\it{e}}(r(z_{t}),D_{w}^{\operatorname{imp}}(\lambda)^{\top}\,\partial_{1}\bar{E}(z_{t}))^{2}]+(1+\lVert D_{w}^{\operatorname{imp}}(\lambda)\rVert_{\sup})^{2}\Big{(}C^{2}_{E,\lambda}\Delta^{2}_{t}+\frac{\sigma_{E,1}+\sigma_{E,2}}{J_{1}}\Big{)}\right).$ (31)
We also note that $\lVert D_{w}^{\operatorname{imp}}(\lambda)\rVert_{\sup}\leq\lVert D_{\Phi,2}(w(\lambda),\lambda)\rVert_{\sup}/(1-q)$ and that
$\tilde{\mathbb{E}}\big{[}\lVert\partial_{1}\bar{E}(z_{t})\rVert^{2}\big{]}\leq 2\tilde{\mathbb{E}}\big{[}\lVert\partial_{1}\bar{E}(z_{t})-\partial_{1}E(z_{t})\rVert^{2}\big{]}+2\lVert D_{E,1}(z_{t})\rVert^{2}_{\sup}\leq\frac{2\sigma_{E,1}}{J_{1}}+2B_{E,\lambda}.$ (32)
Therefore, taking the total expectation in (E.2) and applying Theorem 5.5 with $y=\partial_{1}\bar{E}(z_{t})$ we get